a spark plug in an aoutomobile engine consists of two metal conductors that are separated by a distance of 0.5mm. When an electric spark jumps between them, the magnitude of the electric field is 4.0x107V/m. What is the magnitude of the potential difference V between the conductors?


1
Expert's answer
2021-04-29T10:45:49-0400

The potential difference (in case of uniform electric field) is given as follows:


"\Delta \varphi = Ed"

where "E = 4\times 10^{7}V/m" is the electric field, and "d = 0.5mm = 0.0005m" is the distance between two conductors. Thus, obtain:

"\Delta \varphi =4\times 10^{7}V/m\cdot 0.0005m = 2\times 10^4V"

Answer. "2\times 10^4V".